OpenStudy (anonymous):

Change (3/2 , -3√ ̅3/2) to polar coordinates.

OpenStudy (irishboy123):

(3/2 , -3√ ̅3/2) = 3/2 (1 , -√ ̅3) that should make it easier

OpenStudy (anonymous):

Okay, so then what should I do next?

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (irishboy123):

what do you think the angle is? it's a common one.

OpenStudy (anonymous):

I think I got the answer. But just to be sure, could you check? I got \[\left( 3,\frac{ \pi }{ 3 } \right)\]

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (irishboy123):

it's minus root (3)

OpenStudy (irishboy123):